A point charge with a charge q_1 = 3.70 mu C is he
Answer #1

work done = change in potential energy

initial potential energy = k * q1 * q2 / r

initial potential energy = 9 * 10^9 * 3.7 * 10^-6 * -4.3 * 10^-6 / 0.13

initial potential energy = -1.01 J

final potential energy = 9 * 10^9 * 3.7 * 10^-6 * -4.3 * 10^-6 / sqrt(0.29^2 + 0.28^2)

final potential energy = -0.355 J

change in potential energy = -0.355 - (-1.01)

change in potential energy = 0.655 J

so,

work done by electric force on q2 = 0.655 J